# Examining Caregiving Roles to Differentiate the Effects of Using a Mobile App for Community Oversight for Privacy and Security

### v1

@@ 6. Discussion
Another potential limitation is the risk of privacy breaches if a community member’s phone is lost or stolen, as this could expose the personal information of the entire community. Additionally, in large communities that include members who may not be close, even though were initially trusted, could potentially become malicious. Future research should investigate the possibility of allowing more granular control over the app-hiding features, enabling community members to selectively turn off sharing with specific individuals. Furthermore, caregivers’ advice and privacy behaviors may sometimes be misleading. In such cases, collaborative approaches need to identify deserving candidates to label as caregivers. Factors such as community members’ power use, the validity of their advice, and their privacy practices could be important considerations for future studies. While we allowed all community members to hide any of their installed apps as they preferred, certain mobile apps carry different security and privacy risks (Reardon et al., 2019; Tandon et al., 2022). For example, caregivees were more likely to hide social media and shopping apps, which might have limited the feedback received for those apps. Future research should explore assigning risk-based weights to the apps installed, making users more aware of the risks associated with hiding certain apps and encouraging them to consider sharing to receive effective oversight. Finally, in this study, we did not measure power use in the post study survey as we did not expect the CO-oPS app intervention to alter individual power usage. In retrospect, heightened self-efficacy might lead to increased power usage. Hence, we recommend that future studies on privacy interventions assess this factor both pre and post-intervention to gauge potential changes in power use.
@@ Acknowledgements.
We thank the individuals who participated in our study. We also acknowledge the contributions of Nikko Osaka, Anoosh Hari, and Ricardo Mangandi, in the CO-oPS app development. Our research was supported by the U.S. National Science Foundation under grants CNS-1814068, CNS-1814110, and CNS-2326901. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the U.S. National Science Foundation.

### v2

@@ 6. Discussion
Another potential limitation is the risk of privacy breaches if a community member’s phone is lost or stolen, as this could expose the personal information of the entire community. Additionally, in large communities that include members who may not be close, even though they were initially trusted, they could potentially become malicious. Future research should investigate the possibility of allowing more granular control over the app-hiding features, enabling community members to selectively turn off sharing with specific individuals. Furthermore, caregivers’ advice and privacy behaviors may sometimes be misleading. In such cases, collaborative approaches need to identify deserving candidates to label as caregivers. Factors such as community members’ power use, the validity of their advice, and their privacy practices could be important considerations for future studies. While we allowed all community members to hide any of their installed apps as they preferred, certain mobile apps carry different security and privacy risks (Reardon et al., 2019; Tandon et al., 2022). For example, caregivees were more likely to hide social media and shopping apps, which might have limited the feedback received for those apps. Future research should explore assigning risk-based weights to the apps installed, making users more aware of the risks associated with hiding certain apps and encouraging them to consider sharing to receive effective oversight. Finally, in this study, we did not measure power use in the post study survey as we did not expect the CO-oPS app intervention to alter individual power usage. In retrospect, heightened self-efficacy might lead to increased power usage. Hence, we recommend that future studies on privacy interventions assess this factor both pre and post-intervention to gauge potential changes in power use.
@@ Acknowledgements.
We thank the individuals who participated in our study. We also acknowledge the contributions of Nazmus Sakib Miazi, Nikko Osaka, Anoosh Hari, Madiha Tabassum, and Ricardo Mangandi, in the CO-oPS app development. Our research was supported by the U.S. National Science Foundation under grants CNS-1814068, CNS-1814110, and CNS-2326901. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the U.S. National Science Foundation.